OpenStudy (anonymous):

Use inductive reasoning to predict the next line in the pattern. 18 + 81 = 99 19 + 91 = 110 A. 101 + 20 = 121 B. 20 + 101 = 121 C. 88 + 33 = 121 D. 20 + 101 = 162

OpenStudy (anonymous):

I would go with C

OpenStudy (texaschic101):

18 + 81 = 99 19 + 91 = 110 20 + 101 = 121